40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9082     Entry time: Wed Aug 28 07:33:51 2013
Author: manasa 
Type: Update 
Category: LSC 
Subject: MICH locking 

I wanted to measure the OLTF of MICH.

What I did:
1. Ran LSC offsets script to zero all the offsets.

2. Restored the IFO configure settings for locking Michelson (locked on AS55Q).

3. MICH wouldn't lock on these settings.

4. The MICH servo was hitting its limits (10000 counts). I checked the filter module. After a little bit of looking into things, I disabled FM3 (0,0:5,5), FM4 (1:10) and FM7 (1:5). FM3 and FM7 were filter modules that were switched ON at the trigger. I set these to manual. Enabling any of the filters (FM3, FM4, FM7) caused MICH to lose lock.

5. MICH gain was changed from -20 to -30. MICH locked with ASDC suppressed to 0.01 counts. I looked at the power spectrum of C1:LSC-MICH_OUT on dtt. //edit: Manasa// The plot (uncalibrated) now shows MICH_OUT power spectrum with MICH PSL shutter closed, free-running MICH and loop-enabled MICH.

6. I then wanted to measure the OLTF of MICH using dtt. A channels were set to C1:LSC-MICH_IN1/C1:LSC-MICH_IN2 and excitation given through C1:LSC-MICH_EXC. But I have not been able to get any good coherence for the measurement as yet.

Attachment 1: MICH.pdf  37 kB  Uploaded Wed Aug 28 11:02:44 2013  | Hide | Hide all
MICH.pdf
ELOG V3.1.3-